OpenStudy (anonymous):

Express sin theta in terms of cos theta....HELP PLEASE, Medal to anyone who can help!

OpenStudy (anonymous):

\[\sin^2(\theta)+\cos^2(\theta) = 1\]Solce for cos theta

OpenStudy (anonymous):

*solve

OpenStudy (anonymous):

And that's it? @Gralwyn